OpenStudy (anonymous):

what is the slope of the line defined by y=8x-7

DivineSolar (divinesolar):

Answer choices. Because im gonna walk you through this.

OpenStudy (anonymous):

what are you tryna say here i since a bit of passive aggressiveness

DivineSolar (divinesolar):

Uh? No...

DivineSolar (divinesolar):

Im following the rules.

DivineSolar (divinesolar):

I asked what are the choices.

DivineSolar (divinesolar):

@Twiztidsplifff Answer choices?

OpenStudy (leader):

y=mx+b y=8x-7 what ever is m is your slope @Twiztidsplifff

OpenStudy (anonymous):

that was the question in my book thats all i know

OpenStudy (leader):

i didn't give it to them @DivineSolar I am showing them how you do it...

OpenStudy (anonymous):

wow divine gtfo

DivineSolar (divinesolar):

Not being mean he didnt walk you through it.

DivineSolar (divinesolar):

You have to make them do some work as well.

OpenStudy (leader):

I just showed them how you would find out the slope...

OpenStudy (anonymous):

when you f with my life by f ing with my school then i get mad leave go get a life

OpenStudy (leader):

I didn't give them the answer.

OpenStudy (anonymous):

so your response to him doing that is reporting us ?

DivineSolar (divinesolar):

You arent a moderator bud. You would have purple.

OpenStudy (leader):

@DivineSolar You explain it.. Lets see how you would do it.

OpenStudy (anonymous):

yea we good leader no worrys bro

DivineSolar (divinesolar):

Not you lol @Twiztidsplifff

OpenStudy (leader):

If you are so helpful, you would have a higher SS..

DivineSolar (divinesolar):

Let's solve for x. y=8x−7 Step 1: Flip the equation. 8x−7=y Step 2: Add 7 to both sides. 8x−7+7=y+7 8x=y+7 Step 3: Divide both sides by 8. 8x 8 = y+7 8 x= 1 8 y+ 7 8 Answer: x= 1 8 y+ 7 8 Refresh the page i copied my original answer for this.

OpenStudy (anonymous):

yea divine make yourself usefull

OpenStudy (leader):

I joined two days ago and I have as much as fans and almost the same SS...

OpenStudy (anonymous):

well that was easy wasnt it thank you

DivineSolar (divinesolar):

I honestly dont believe you leader but still i did a better explanation than him. ^_^

OpenStudy (leader):

ANd your reply is not even readable... Fail.

DivineSolar (divinesolar):

Wouldnt you like a graph as well X3

DivineSolar (divinesolar):

Sure beacuse you didnt refresh the page as i said to

OpenStudy (leader):

You didn't answer their question.. YOu don't even know math.

DivineSolar (divinesolar):

Aw it doesnt matter im still a fan of urs X)

DivineSolar (divinesolar):

And yes i did lol, slope is in the answer. 1/8

OpenStudy (leader):

Actually slope is 8...

OpenStudy (leader):

Is this readable @DivineSolar http://prntscr.com/55zam2

DivineSolar (divinesolar):

Yes its fine for me XD

DivineSolar (divinesolar):

I dont see anything wrong with it?

DivineSolar (divinesolar):

There are no question mark things for me.

DivineSolar (divinesolar):

If thats what you are implieng.

DivineSolar (divinesolar):

@leader we could both be wrong X3

DivineSolar (divinesolar):

Btw slope is rise over run so. Your wrong X3

OpenStudy (anonymous):

wow